Archer Contracting repaved 180 miles of 3-lane roadway with a crew of 10 employees. This crew worked 25 days and used $9,000 worth of paving material. Nearby, Bronson Construction repaved 150 miles of 4-lane roadway working 20 days with a crew of 15 employees using $10,000 of paving material. Both Archer Contracting and Bronson Construction rented the same paving equipment to complete their respective assignments, costing each $2,000 a day.

a. Calculate the labor productivity of both Archer Contracting and Bronson Construction. Who is more productive and what is the percent difference?

b. Assume that both Archer Contracting and Bronson Construction pay their employees $21 an hour for 8 hours per day. Calculate the multifactor productivity of each firm. Now who is more productive and what is the percent difference?
Answer #1

Labor productivity = Total miles repaved/(Total crew members*Total working days)

a) Labor productivity:

Archer Contracting = 180/(10*25) = 0.72 miles per man day

Bronson Construction = 150/(15*20) = 0.5 miles per man day

So, Archer contracting is better

b)

Multifactor productivity = Total miles repaved/(Labor cost + Material cost + Equipment cost)

Labor cost = Number of crew*Working days*Hours per day*Hourly wage; Hours per day = 8; Hourly wage = 21$

Equipment cost = Working days*Equipment cost per day

Archer contracting = 180/(10*25*8*21 + 9000 + 25*2000) = 0.00178 miles per dollar input

Bronson Construction = 150/(15*20*8*21 + 10000 + 20*2000) = 0.00149 miles per dollar input

So, Archer contracting is better